OpenStudy (anonymous):

Distribute 6(3x+5)-4x=35

OpenStudy (igreen):

6(3x + 5) - 4x = 35 Distribute: 18x + 30 - 4x = 35 Subtract 30 to both sides: 18x - 4x = 5 Simplify: 14x = 5 Divide 14 to both sides: 0.36

OpenStudy (igreen):

Your answer is going to be: 18x + 30 - 4x = 35 If you just distribute.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @RideRANK22